# FunDiff: Diffusion Models over Function Spaces for Physics-Informed Generative Modeling

### v1

@@ 4 Results
The DiT follows a similar optimization strategy but operates with larger batches of 128 samples over $10^{5}$ iterations. For training the conditioned diffusion models, we generate conditions by randomly downsampling input functions and encoding them through the pretrained FAE encoder. Detailed hyperparameters for all experiments are provided in Appendix C.
4.1 Learning 1D damped sinusoidal functions
$f(x)=A\exp(-\gamma x)\sin(\omega x)+b,\quad x\in[0,1],$ (4.1)
We employ an FAE with a 6-layer encoder and a 2-layer decoder, each with a hidden dimension of 256 and 8 attention heads. The FAEs are trained on datasets of varying sizes, with function samples discretized at a resolution of 128. We training FAE by minimizing the loss using the Adam optimizer for $5\times 10^{4}$ iterations. The learning rate follows a warmup and exponential decay schedule: it linearly increases from 0 to $10^{-3}$ over the first 1,000 steps, then decays exponentially every 1,000 steps with a decay rate of 0.9. Following the FAE training, we train the latent diffusion model by minimizing Eq. (3.2) using an identical learning rate schedule.
For evaluation, we generate 8,192 sample functions using our inference procedure with 100 denoising steps. We assess the quality of these generated samples through parameter estimation, leveraging the analytical form of damped sinusoidal functions. Specifically, for each generated sample, we perform least-squares fitting using Eq. (4.1) to obtain the inferred parameters and reconstruct the corresponding signal. This approach enables a rigorous quantitative assessment of our model’s ability to capture the underlying function space through two metrics: the comparison between reconstructed and generated signals, and the analysis of inferred parameter distributions against the target uniform distributions.
We provide a comprehensive evaluation of our model’s generative performance across varying training dataset sizes (Fig. 2). The top row illustrates the worst-case generated samples alongside their reconstructions via parameter estimation. As the number of training samples increases, the generation quality improves significantly. The bottom row analyzes the parameter distributions, revealing the evolution of our model’s capacity to capture the underlying function space. With limited training data (16 or 32 samples), the estimated parameter distributions exhibit narrower ranges or deviate significantly from the target uniform distributions, indicating insufficient coverage of the parameter space. As the training dataset size increases, the estimated parameter distributions progressively converge to the target uniform distributions, with the model trained on 256 samples achieving notably better alignment with the true distributions. This improvement is quantitatively validated by the decreasing mean squared error (MSE) between the generated samples and their reconstructions. These empirical results strongly support our theoretical analysis in Section 3.
Figure 2: Damped sinusoidal functions. Analysis of generative quality across models trained with different numbers of samples ($n=16,32,64,128,256$). (Top) Comparison between the worst-case generated samples (blue dots) and their reconstructed signals (red dashed lines) obtained through parameter estimation. (Bottom) The first four panels show the estimated parameter distributions for amplitude ($A$), damping coefficient ($\gamma$), angular frequency ($\omega$), and vertical shift ($b$). Red dashed lines indicate the true parameter bounds from the underlying uniform distributions. The rightmost panel displays the MSE of the reconstructed signal. The results demonstrate that models trained with more samples generate functions that better align with the true distributions and achieve lower reconstruction error.
4.2 Kolmogorov flow
In this example, we use FunDiff to generate divergence-free velocity fields, a fundamental requirement for ensuring the physical consistency of incompressible flow simulations in computational fluid dynamics. We consider two scenarios: unconditional flow generation and flow reconstruction from partial observations. To evaluate our approach, we use the two-dimensional Kolmogorov flow, a well-established benchmark for analyzing transition to turbulence and the stability of incompressible fluid systems. The governing equations are formulated in the vorticity-velocity representation as follows:
$\frac{\partial\omega(x,t)}{\partial t}+\mathbf{u}(x,t)\cdot\nabla\omega(x,t)$ $=\frac{1}{\operatorname{Re}}\nabla^{2}\omega(x,t)+f(x),\quad x\in(0,2\pi)^{2},t\in(0,T],$
$\nabla\cdot\mathbf{u}(x,t)$ $=0,\quad x\in(0,2\pi)^{2},t\in(0,T],$
$\omega(x,0)$ $=\omega_{0}(x),\quad x\in(0,2\pi)^{2},$
where $\omega$ is the vorticity, $\mathbf{u}=(u,v)$ is the velocity field, and Re is the Reynolds number, set to 1000. The forcing term $f(x)$ drives the flow, and $x=\left[x_{1},x_{2}\right]$. Periodic boundary conditions are applied. The forcing term is defined as
$f(x)=-4\cos\left(4x_{2}\right)-0.1\omega(x,t),$
where the additional drag force term $0.1\omega(x,t)$ prevents energy accumulation at large scales.
$\left[x,y,\mathcal{E}_{\theta}(f)\right]\xrightarrow{\mathcal{D}_{\theta}}\psi(\gamma(x,y)),$
$\gamma(x,y)=[\cos(x),\sin(x),\cos(y),\sin(y)].$
This embedding guarantees that the output functions have a periodicity of $2\pi$ within $[0,2\pi]^{2}$. An incompressible velocity field is then computed as $(u,v)=(\partial\psi/\partial y,-\partial\psi/\partial x)$.
4.2.1 Unconditional flow generation
Following the training methodology in Section 3.3, we present our results in Fig. 3a, which displays representative generated samples along with their corresponding divergence fields, computed via automatic differentiation. The predicted divergence values are negligible, confirming the effectiveness of our approach. We also compare the spectral energy distributions between the training dataset and 1,000 generated samples (Fig. 3b). The spectra show excellent agreement up to wavenumbers of $\mathcal{O}(10^{2})$. However, the model leads to some artificial high-frequency components at larger wavenumbers.
This unconstrained approach results in significantly larger divergence values (Fig. 7). Additionally, the spectral energy shows some discrepancies compared to the model constrained with the exact divergence-free condition.
4.2.2 Flow reconstruction from partial observations
Furthermore, we benchmark our approach against several state-of-the-art neural operators and PDE surrogate models, including FNO [14], ViT [8], and UNet [60]. Detailed implementation specifications for these baseline models are provided in Appendix C.1. Fig. 3c presents reconstructed flow fields from various models, where the input consists of 64 × 64 resolution velocity fields corrupted by 20% Gaussian noise. The results demonstrate that our model uniquely achieves both smooth and accurate reconstructions. Additional visualizations of our model’s predictions, with and without the divergence-free condition imposed in FAE decoder, are shown in Figs. 8 and 9, respectively. The quantitative comparison using relative $L^{2}$ errors across different downsampling factors is presented in Fig. 3e. These results confirm that our FunDiff model maintains consistent performance across all tested resolutions while exhibiting strong resilience to noise.
4.3 Burgers’ equation
As our next example, we demonstrate the effectiveness and flexibility of our framework to generate physically consistent samples through incorporating physical laws as soft constraints in the training of FAE. To this end, we consider the Burgers’ equation, which serves as a fundamental prototype for conservation laws that can develop discontinuities. The governing equation is given by
$\frac{du}{dt}+u\frac{du}{dx}-\nu\frac{d^{2}u}{dx^{2}}=0,\quad(x,t)\in(0,1)\times(0,1],$ (4.2)
$u(x,0)=a(x),\quad x\in(0,1)$ (4.3)
with periodic boundary conditions, where $\nu=0.001$ denotes the viscosity coefficient, and the initial condition $a(x)$ is sampled from a Gaussian random field (GRF) satisfying the periodic boundary conditions.
Let $u_{\theta}$ denote the FAE output. We define the PDE residual as
The FAE is trained by minimizing a composite loss
$\mathcal{L}(\mathbf{\theta}):=\underbrace{w_{data}\frac{1}{NM}\sum_{i=1}^{N}\sum_{j=1}^{M}\left|{u}_{\mathbf{\theta}}^{i}(x_{j},t_{j})-{u}^{i}(x_{j},t_{j})\right|^{2}}_{\mathcal{L}_{data}(\mathbf{\theta})}+\underbrace{w_{physics}\frac{1}{NM}\sum_{i=1}^{N}\sum_{j=1}^{M}\left|\mathcal{R}[{u}_{{\theta}}^{i}](x_{j},t_{j})\right|^{2}}_{\mathcal{L}_{physics}(\mathbf{\theta})}.$
Similar to the example of Kolmorgorv flow, we evaluate our model against established baseline methods. We show three examples in Fig. 4a. We can observe that FNO, ViT, and UNet predict blurry discontinuous predictions. In contrast, our FunDiff generates sharp and smooth reconstructions, which are in excellent agreement with the reference numerical solutions. Fig. 4b presents a quantitative comparison of model performance across inputs with varying downsampling factors and 20% Gaussian noise corruption. Notably, FunDiff outperforms all baselines. Additional examples of our model’s predictions are shown in Fig. 10. Their physical consistency is confirmed by low PDE residuals, remaining on the order of $\mathcal{O}(10^{-2})$ across the entire domain. We also systematically examine the effect of noise and data sparsity (Fig. 4c). As expected, reconstruction error increases with higher noise levels, but FunDiff demonstrates strong robustness to data sparsity, maintaining stable performance even under aggressive downsampling (factors up to 5).
To rigorously evaluate the impact of physical constraints, we conduct an ablation study by training the function encoder without the physics loss term $\mathcal{L}_{physics}$. The absence of PDE-based regularization leads to significantly higher predictive errors and PDE residuals in the generated solutions (Fig. 4d). This degradation in physical consistency is further visualized through the generated samples in Fig. 11, underscoring the crucial role of physics-informed training in our framework.
4.4 Linear elasticity
Symmetries play an important role in physics, providing a framework for understanding conservation laws and invariances that govern natural phenomena. This example aims to demonstrate how to embed prior symmetries into generative models. To this end, we consider linear elasticity in computational micromechanics, which involves solving the Lippmann-Schwinger equation [61] in a 2D periodic representative volume element (RVE) domain $\Omega\subset\mathbb{R}^{2}$ with spatial coordinates $\mathbf{x}=(x,y)$ under plane strain conditions. The problem is formulated as
$\begin{dcases*}\varepsilon_{ij}(\mathbf{x})+\int_{\Omega}G_{ijkl}^{(0)}\left(\mathbf{x},\mathbf{x}^{\prime}\right):[\sigma_{ij}\left(\mathbf{x}^{\prime}\right)-C_{ijkl}^{0}:\varepsilon_{kl}\left(\mathbf{x}^{\prime}\right)]\mathrm{d}\mathbf{x}^{\prime}-\bar{\varepsilon}_{ij}=0,\quad\mathbf{x}\in\Omega,\\ \sigma_{ij}(\mathbf{x})=C_{ijkl}(\mathbf{x}):\varepsilon_{kl}(\mathbf{x}).\\ \end{dcases*}$ (4.4)
These relations represent balance and constitutive equations, respectively, which are written in a component-wise manner ($1\leq i,j,k,l\leq 2$) using Voigt notation. $\mathbf{x}$ and $\mathbf{x}^{\prime}$ denote a pair of target and source points inside the RVE domain $\Omega$. $\sigma_{ij},\varepsilon_{ij}$ and $C_{ijkl}$ denote stress, strain, and fourth-order elastic tensor, respectively. $C_{ijkl}(\mathbf{x})$ is expressed in terms of local Young’s modulus $E(\mathbf{x})$ and Poisson ratio $\nu(\mathbf{x})$ as
where $\delta_{ij}$ denotes Kronecker delta function. $G_{ijkl}^{(0)}$ denotes Green’s operator corresponding to the homogeneous reference material with elastic tensor $C_{ijkl}^{0}$ [62]. $\bar{\varepsilon}_{ij}$ denotes far field strain used to apply the strain-controlled boundary condition, equal to the average strain over the RVE domain $\Omega$. In this example, we model the point-wise strain concentration tensor $A_{ijkl}(\mathbf{x})\in\mathbb{R}^{3\times 3}$, which linearly maps $\bar{\varepsilon}_{ij}$ to $\varepsilon_{ij}(\mathbf{x})$ independent of loading conditions:
$A_{ijkl}(\mathbf{x})$ can be computed by triply solving Eq. (4.4) subject to three orthogonal unit far field strain $\overline{\varepsilon}_{ij}$ as
$\overline{\varepsilon}_{ij}=\begin{Bmatrix}[1,0,0]^{T},[0,1,0]^{T},[0,0,1]^{T}\end{Bmatrix}.$ (4.5)
Due to intrinsic symmetry of the RVE configuration, the resulting strain concentration tensor field $A_{ijkl}(\mathbf{x})$ exhibits a symmetric pattern with respect to both $x=\frac{1}{2}$ and $y=\frac{1}{2}$. We modify the forward pass of the decoder to enforce these constraints. Let $\mathbf{z}$ denote the FAE encoder output latents. The symmetry-preserved decoder $\hat{\mathcal{D}}_{\theta}$ is then constructed as
Fig. 5a presents reconstructed samples generated by our FunDiff model from a coarse strain concentration field corrupted by 10% Gaussian noise. The results show that our model preserves underlying symmetries and accurately resolves sharp material interfaces. To assess robustness against different noise types, we report test errors in Fig. 5b. The consistently low errors across all noise configurations highlight the framework’s reliability in practical applications.
To further demonstrate the critical role of the continuous decoder in our framework, we implement a baseline using a discrete autoencoder with a Vision Transformer encoder and an MLP decoder projecting to the target pixel space. We train a diffusion transformer under identical conditions, with representative generated samples shown in Fig. 12. Unlike our approach, the baseline fails to preserve symmetry or enforce physical continuity in strain fields, instead producing piecewise discontinuous predictions.
4.5 Turbulence mass transfer
As our last example, we examine the model capacity of inferring non-intrusive quantities such as pressure in turbulent problems. Specifically, we consider turbulent mass transfer in a rectangular channel. This phenomenon plays a critical role in chemical processing, directly influencing reaction kinetics, conversion efficiency, and product purity. Our study focuses on a computational domain $\Omega=\{(x,y)\mid 0.006<x<0.018,0<y<0.01\}$. Within this domain, we are interested in inferring the obstacle topology and pressure field given the velocity measurements.
$\frac{\partial(\rho u_{i})}{\partial x_{i}}=0,\quad$ (Continuity equation)
$u_{i}\frac{\partial(\rho u_{j})}{\partial x_{i}}=-\frac{\partial P}{\partial x_{j}}+\frac{\partial}{\partial x_{i}}\left[\left(\mu_{m}+\mu_{t}\right)\frac{\partial u_{j}}{\partial x_{i}}\right],\quad$ (Momentum conservation equation)
$u_{i}\frac{\partial(\rho C)}{\partial x_{i}}=\frac{\partial}{\partial x_{i}}\left[\rho\left(D_{m}+D_{t}\right)\frac{\partial C}{\partial x_{i}}\right],\quad$ (Species conservation equation)
where $p,u_{i},x_{i}$ are the pressure, velocity, and location respectively; the subscript $i$ represents the coordinate direction which can take $x$ and $y$ for the two-dimensional system. Here, $C$ is the time-averaged concentration; $\rho,\mu_{m}$ and $D_{m}$ are density, laminar viscosity, and laminar mass diffusivity, which are physical property constants and fixed to $1000\mathrm{~{}kg}/\mathrm{m}^{3},1.003\times 10^{-6}\mathrm{~{}m}^{2}/\mathrm{s},1.003\times 10^{-8}\mathrm{~{}m}^{2}/\mathrm{s}$, respectively. $\mu_{t}$ and $D_{t}$ are the turbulent viscosity and turbulent mass diffusivity, respectively.
Representative generated samples (Fig. 6a) validate the model’s ability to infer pressure fields and reconstruct obstacle geometries despite heavily corrupted input velocity measurements. The reconstructions closely align with numerical estimations. Fig. 6b presents the test error across various downsampling factors, noise levels, and noise types. The results indicate that our model remains robust even under severe noise corruption and low-resolution inputs. Our results demonstrate the model’s effectiveness in preserving critical flow features and geometric details, and highlight the framework’s potential for real-world applications where high-fidelity reconstructions are required from sparse or noisy measurements.
5 Discussion
In this work, we propose FunDiff, a flexible and efficient framework for generative modeling in function spaces, addressing fundamental challenges in generating continuous function data while preserving physical constraints. Our primary contributions are twofold: (1) a rigorous theoretical foundation establishing conditions for function distribution approximation via autoencoders, and (2) a novel constrained function autoencoder (FAE) that inherently enforces physical principles. Through extensive experimentation across diverse physical systems, we have demonstrated that our approach successfully generates physically valid samples while maintaining high fidelity to the target distribution.
While our current work lays a strong foundation for physics-informed generative modeling, several promising directions for future research emerge. A natural extension would be to adapt our framework to handle more complex geometries, thereby enabling broader applicability in real-world scientific and engineering problems. Handling complex domains necessitates incorporating geometric inductive biases, such as mesh-based representations, implicit neural surfaces, or graph-based function approximations. Moreover, extending our approach to multiphysics applications represents a critical and exciting frontier. Many engineering and scientific challenges involve the interplay of multiple physical processes, such as fluid-structure interactions, thermo-mechanical coupling, and electromagnetic-thermal phenomena, which require models that can simultaneously capture and reconcile diverse physical constraints. These advancements will significantly expand the scope of generative modeling for physical systems and ultimately contribute to more accurate and efficient simulation-based design across disciplines.
@@ Acknowledgments
LL acknowledges support from the U.S. Department of Energy, Office of Advanced Scientific Computing Research, under Grants No. DE-SC0025593 and No. DE-SC0025592, and the U.S. National Science Foundation under Grant No. DMS-2347833. SW acknowledges support from the NVIDIA Academic Grant Program. We thank Dr. Chenhui Kou for providing the turbulence mass transfer dataset and Dr. Mingyuan Zhou for insightful discussions.
Appendix A Notations
Our notations are summarized in Table 1.
Table 1: Main symbols and notations used in this work.
Notation Description
FunDiff
@@ Appendix A Notations
$W_{1}(\mu,\nu)=\sup_{f\text{~{}is 1-Lipschitz}}\left[\int_{\mathcal{X}}f\,{\mathrm{d}}\mu-\int_{\mathcal{X}}f\,{\mathrm{d}}\nu\right].$
@@
Appendix B Theoretical proofs
@@
B.1 Density estimation error decomposition
In this part, we prove the decomposition of density estimation Wasserstein error in Eq. (3.1). According to the FAE framework, we have $\widehat{P}=G_{\sharp}\hat{p}$, which leads to
$W_{1}(P,\widehat{P})$ $=W_{1}(P,G_{\sharp}\hat{p})\overset{(a)}{\leqslant}W_{1}(P,G_{\sharp}p)+W_{1}(G_{\sharp}p,G_{\sharp}\hat{p})$
$\overset{(b)}{\leqslant}{\mathbb{E}}_{f\sim P}\|f-G\circ E(f)\|_{2}+\mathrm{Lip}(G)\cdot W_{1}(p,\widehat{p}).$
@@ B.1 Density estimation error decomposition
$W_{1}(G_{\sharp}p,G_{\sharp}\hat{p})\leqslant\sup_{\theta,\theta^{\prime}}\frac{\|G(\theta)-G(\theta^{\prime})\|_{2}}{\|\theta-\theta^{\prime}\|_{2}}\cdot W_{1}(p,\hat{p})=\mathrm{Lip}(G)\cdot W_{1}(p,\hat{p}).$
@@
holds for $\forall f\in\mathcal{F}_{s}$ according to the approximation error in Lemma 3.4, which leads to
@@
This lemma concludes that $\mathrm{Lip}(G)\lesssim\sqrt{d/m}$ in Eq. (3.1). Finally, the latent dimension $D=m(2+d)$ and the oracle loss
@@
Appendix C Experimental details
C.1 Baselines
In our experiments, we compare our method against three widely used baselines—Fourier Neural Operator (FNO), Vision Transformer (ViT), and UNet—which represent diverse strategies for capturing function dependencies and complex data representations. We train all baseline models using the same training procedure described in Section 4. For random downsampled inputs, we first interpolate them back to the original resolution before being processed.

### v2

@@ Methods
The DiT follows a similar optimization strategy but operates with larger batches of 128 samples over $10^{5}$ iterations. For training the conditioned diffusion models, we generate conditions by randomly downsampling input functions and encoding them through the pretrained FAE encoder. Detailed hyperparameters for all experiments are provided in Appendix S4.
@@ Acknowledgments
LL acknowledges support from the U.S. Department of Energy, Office of Advanced Scientific Computing Research, under Grants No. DE-SC0025593 and No. DE-SC0025592, and the U.S. National Science Foundation under Grants No. DMS-2347833 and No. DMS-2527294. SW acknowledges support from the NVIDIA Academic Grant Program. We thank Dr. Chenhui Kou for providing the turbulence mass transfer dataset and Dr. Mingyuan Zhou for insightful discussions.
Author contributions statement
S.W., Z.D. and L.L. conceived the project and developed the main methodology. S.W. implemented the core algorithms and conducted the experiments. Z.D. contributed to the theoretical analysis. T.L. and S.S. assisted with numerical simulation. L.L. supervised the project and provided critical revisions. All authors discussed the results and contributed to writing the manuscript.
Appendix S1 Notations
Our notations are summarized in Table S1.
Table S1: Main symbols and notations used in this work.
[…]
@@
$W_{1}(\mu,\nu)=\sup_{f\text{\penalty 10000\ is 1-Lipschitz}}\left[\int_{\mathcal{X}}f\,{\mathrm{d}}\mu-\int_{\mathcal{X}}f\,{\mathrm{d}}\nu\right].$
@@
Appendix S2 Preliminaries
Before introducing our generative models in function spaces, we provide some basic knowledge of diffusion models and its Minimax optimality.
We begin by reviewing foundational concepts of diffusion models in both discrete and continuous settings. Diffusion models form a class of generative frameworks that approximate complex data distributions through a two-phase process: a forward noising process that incrementally corrupts the data, followed by a backward denoising process that reconstructs data from noise.
$q(\mathbf{x}_{1:N}\mid\mathbf{x}_{0})$ $=\prod_{k=1}^{N}q(\mathbf{x}_{k}\mid\mathbf{x}_{k-1}),$ (S1)
$q(\mathbf{x}_{k}\mid\mathbf{x}_{k-1})$ $=\mathcal{N}\bigl(a_{k}\mathbf{x}_{k-1},\,b_{k}^{2}\mathbf{I}\bigr).$ (S2)
$\frac{\mathrm{d}{\mathbf{x}}}{\mathrm{d}t}=\mathbf{v}({\mathbf{x}},t),$
Density estimation and the minimaxity of diffusion models
As a core problem in statistics and machine learning, the purpose of density estimation is to solve the following problem: given $n$ i.i.d. data points $\{x_{i}\}_{i=1}^{n}$ sampled from an unknown distribution $p^{*}$, we want to construct an approximate distribution $\hat{p}$ based on the $n$ given data points. Density estimation arises in diverse applications, from anomaly detection and clustering to modern generative modeling, where having an accurate handle on the underlying distribution is crucial. One broad class of solutions is parametric methods, which assume that the ground truth distribution $p^{*}$ lies within a known family of distributions $\{p_{\theta}:\penalty 10000\ \theta\in\Theta\}$. The density estimation problem then reduces to finding the optimal parametric vector $\hat{\theta}$ via log-likelihood maximization:
$\hat{p}_{h}(x)=\frac{1}{nh^{d}}\sum_{i=1}^{n}K\left(\frac{x-x_{i}}{h}\right).$
Recently, diffusion models have emerged as powerful generative approaches that approximate the data distribution by reversing a forward noise-adding process. Although widely recognized for their empirical success, these models also admit a theoretical interpretation under the lens of minimax optimality for density estimation. In fact, under suitable assumptions on the function space that characterizes the ground truth distribution $p^{*}$, one can show that diffusion-based estimators achieve minimax rates for distribution (or score) recovery. Here we outline the theoretical underpinnings of this minimax perspective by first presenting a lower bound result (Theorem S1) that establishes a fundamental limit for density estimation in the space of all continuous functions $\mathcal{C}({\mathbb{R}}^{d})$, and then provide an upper bound result (Theorem S2) that shows how diffusion-based methods (and related estimators) can achieve this rate, thus elucidating the minimax guarantees behind modern generative modeling frameworks.
Theorem S1 (Lower Bound, Theorem 3 of Ref. [82]).
  • (i)
  • (ii)
Theorem S2 (Upper Bound, Section 3.3 of Ref. [54]).
Let $\alpha\geqslant 0$ denote the smoothness parameter, $d\geqslant 1$ the dimension, and $c>0$ a positive constant. We still consider the class of densities $\mathcal{F}\subseteq\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ defined in Theorem S1. By solving the backward diffusion SDE
${\mathrm{d}}Y_{t}=\hat{s}(Y_{t},T-t){\mathrm{d}}t+{\mathrm{d}}W_{t}$
Appendix S3 Theoretical proofs
@@
S3.1 Density estimation error decomposition
In this part, we prove the decomposition of density estimation Wasserstein error in (1). According to the FAE framework, we have $\widehat{P}=G_{\sharp}\hat{p}$, which leads to
[…]
@@
[…]
@@
holds for $\forall f\in\mathcal{F}_{s}$ according to the approximation error in Lemma 4, which leads to
@@
This lemma concludes that $\mathrm{Lip}(G)\lesssim\sqrt{d/m}$ in (1). Finally, the latent dimension $D=m(2+d)$ and the oracle loss
@@
S3.4 Lower bound: Proof of Theorem 6
Step 1: Bi-Lipschitz embedding of a Euclidean ball into RKHS Ball $\mathcal{B}$
For a fixed dimension $d\geqslant 1$, define the following linear map $T_{d}:{\mathbb{R}}^{d}\rightarrow\mathcal{H}$ by
$S_{d}(t):=\sum_{i=1}^{d}\sqrt{\mu_{i}}\cdot t_{i}\phi_{i},$
where $\{\mu_{i}\}$ and $\{\phi_{i}\}$ are the eigenvalues and eigen-functions of RKHS $\mathcal{H}$, respectively. For vector $t\in{\mathbb{R}}^{d}$, we have
$\left\|T_{d}(t)\right\|_{\mathcal{H}}^{2}=\sum_{i=1}^{d}\frac{(\sqrt{\mu_{i}}\cdot t_{i})^{2}}{\mu_{i}}=\sum_{i=1}^{d}t_{i}^{2}=\|t\|_{2}^{2}.$
Hence $T_{d}:t\in B_{d}(1)\mapsto T_{d}(t)\in\mathcal{B}$, where $B_{d}(1):=\{t\in{\mathbb{R}}^{d}\penalty 10000\ :\penalty 10000\ \|t\|_{2}\leqslant 1\}$ is a $d$-dimensional Euclidean ball. Moreover, for all $s,t\in{\mathbb{R}}^{d}$, we have
$\|T_{d}(t)-T_{d}(s)\|^{2}_{L_{2}(\nu)}=\sum_{i=1}^{d}\mu_{i}(t_{i}-s_{i})^{2}\penalty 10000\ \penalty 10000\ \Rightarrow\penalty 10000\ \penalty 10000\ \sqrt{\min_{i\in[d]}\mu_{i}}\cdot\|t-s\|_{2}\leqslant\|T_{d}(t)-T_{d}(s)\|_{L_{2}(\nu)}\leqslant\sqrt{\max_{i\in[d]}\mu_{i}}\cdot\|t-s\|_{2}.$
Since $\mu_{1}\asymp 1,\mu_{d}\asymp d^{-2\beta}$, the lower Lipschitz factor is
$L_{d}:=\sqrt{\mu_{d}}\geqslant\sqrt{c}\cdot d^{-\beta}.$
Now, we introduce the pushforward class. Let $\mathcal{Q}_{d}$ be the set of probability measures supported on the $d$-dimensional unit Euclidean ball $B_{d}(1)$. Consider its image
$\mathcal{P}_{d}:=\{T_{d}\#Q\penalty 10000\ :\penalty 10000\ Q\in\mathcal{Q}_{d}\}\subseteq\{P\penalty 10000\ :\penalty 10000\ \mathrm{supp}(P)\subseteq\mathcal{B}\}.$
Step 2: Transferring Wasserstein distance through the embedding
For any $Q,Q^{\prime}\in\mathcal{Q}_{d}$, by the definition of $W_{1}$ and the lower Lipschitz bound,
$W_{1}\left(T_{d}\#Q,T_{d}\#Q^{\prime}\right)\geqslant L_{d}\cdot W_{1}^{{\mathbb{R}}^{d}}(Q,Q^{\prime}),$
where $W_{1}^{{\mathbb{R}}^{d}}$ is the Wasserstein-1 metric in ${\mathbb{R}}^{d}$ with Euclidean ground norm. Therefore, we can apply estimator pullback technique and use the statistical lower bound for density estimation inside $d$-dimensional unit ball (Euclidean space) to cap the statistical lower bound for density estimation inside an RKHS ball (function space).
Given any estimator $\widehat{P}_{n}=\widehat{P}_{n}(X_{1:n})$ taking values in probability measures on $\mathcal{B}$, define a corresponding ${\mathbb{R}}^{d}$-estimator $\widehat{Q}_{n}$ on data $Y_{i}\in B_{d}(1)$ by
$\widehat{Q}_{n}(Y_{1:n})\in\arg\min_{Q\in\mathcal{Q}_{d}}W_{1}\left(T_{d}\#Q,\widehat{P}_{n}\left(T_{d}(Y_{1}),T_{d}(Y_{2}),\ldots,T_{d}(Y_{n})\right)\right).$
By the optimality of the argmin and the inequality above, we can conclude that
$W_{1}\left(T_{d}\#Q,\widehat{P}_{n}(T_{d}(Y_{1:n}))\right)\geqslant L_{d}\cdot W_{1}^{{\mathbb{R}}^{d}}(Q,\widehat{Q}_{n}(Y_{1:n})).$
Taking expectation under $Y_{i}\sim Q$ and taking supremum over $Q\in\mathcal{Q}_{d}$, we obtain the risk domination
$\sup_{P\in\mathcal{P}_{d}}{\mathbb{E}}\left[W_{1}(P,\widehat{P}_{n})\right]\geqslant L_{d}\cdot\sup_{Q\in\mathcal{Q}_{d}}{\mathbb{E}}\left[W_{1}^{{\mathbb{R}}^{d}}(Q,\widehat{Q}_{n})\right].$
Finally, we take the infimum over all original estimators $\widehat{P}_{n}$ (which induces taking infimum over all such $\widehat{Q}_{n}$) yields
$\inf_{\widehat{P}_{n}}\sup_{P\in\mathcal{P}_{d}}{\mathbb{E}}\left[W_{1}(P,\widehat{P}_{n})\right]\geqslant L_{d}\cdot\left(\inf_{\widehat{Q}_{n}}\sup_{Q\in\mathcal{Q}_{d}}{\mathbb{E}}[W_{1}^{{\mathbb{R}}^{d}}(Q,\widehat{Q}_{n})]\right):=R_{n}^{(d)}.$
Step 3: A $d$-dimensional Wasserstein lower bound on the Euclidean ball
For each fixed $d\geqslant 1$, a standard conclusion of the classical curse-of-dimensionality for Wasserstein estimation tells us that, on a bounded set of ${\mathbb{R}}^{d}$, the minimax $W_{1}$ risk scales as $R_{n}^{(d)}\gtrsim n^{-1/d}$. Therefore,
$\inf_{\widehat{P}_{n}}\sup_{P\in\mathcal{P}_{d}}{\mathbb{E}}\left[W_{1}(P,\widehat{P}_{n})\right]\geqslant L_{d}\cdot R_{n}^{(d)}\gtrsim d^{-\beta}\cdot n^{-1/d}.$
Step 4: Optimize the bound over dimension $d$
For any $d\geqslant 1$, since $\mathcal{P}_{d}\subset\{P\penalty 10000\ :\penalty 10000\ \mathrm{supp}(P)\subseteq\mathcal{B}\}$, we have
$\mathcal{R}_{n}:=\inf_{\widehat{P}_{n}}\sup_{P:\penalty 10000\ \mathrm{supp}(P)\subseteq\mathcal{B}}\geqslant\inf_{\widehat{P}_{n}}\sup_{P\in\mathcal{P}_{d}}{\mathbb{E}}\left[W_{1}(P,\widehat{P}_{n})\right]\gtrsim d^{-\beta}\cdot n^{-1/d}.$
Now we choose an optimal dimension $d$ to maximize $d^{-\beta}\cdot n^{-1/d}$. By basic calculus, it is obvious that the maximizer $d^{*}=\frac{\log n}{\beta}$, where $d^{-\beta}\cdot n^{-1/d}=(\beta/e)^{\beta}\cdot\left(\frac{1}{\log n}\right)^{\beta}$. To sum up, we have
$\mathcal{R}_{n}:=\inf_{\widehat{P}_{n}}\sup_{P:\penalty 10000\ \mathrm{supp}(P)\subseteq\mathcal{B}}\gtrsim\left(\frac{1}{\log n}\right)^{\beta},$
which comes to our conclusion.
Appendix S4 Experimental details
S4.1 Baselines
In our experiments, all baseline models are trained using the same procedure described in the Results section. Specifically, each model is trained for $10^{5}$ iterations using the AdamW optimizer with a weight decay of $10^{-5}$ and a batch size of 16. The learning rate schedule includes a warm-up of 2,000 steps from zero to the base learning rate. For neural operator baselines, we use a base learning rate of $10^{-3}$ with exponential decay every 2,000 steps at a rate of 0.9. For diffusion-based models, we use a base learning rate of $10^{-4}$ with the same decay schedule, and apply exponential moving average (EMA) updates to stabilize training. For randomly downsampled inputs, we interpolate them back to the original resolution before processing.